1. Let the national income model be Y = C0 + c (Y – 3tY) + I + G + X – M, where C0, I, G, X, and M are given. The government considers increasing G by EUR 1 million. Assuming c = 0.3 and t = 0.15, what would be the expected change in the national output?
1
Expert's answer
2022-01-16T17:43:48-0500

"G0=14"

"Y=C+I0+G0"

"Y= (25+6Y1\/2) 16 + 14"

"Y=55+6Y1\/2"

"Y^2= 55+6Y"

"(Y^2)-6Y-55=0"

"(y-11)(y+5)=0"

"11"

"Y=121"

"121=C+16+14"

"91=C"

"Y*=121"

"C*= 91"

If the marginal propensity to import increases, the multiplier will decrease. Imports, like income taxes, are automatic stablizers. If everyone increases their marginal propensity to save, the Keynesian model predicts that total saving will not increase, and may decline. This is the “paradox of thrift”